OpenStudy (anonymous):

Find the absolute extrema of the function. (Round your answer to three decimal places.) f(x) = xe-x2 on [0,2] Absolute maximum value at x = Absolute minimum value: at x =

OpenStudy (anonymous):

Is that \[x^e - x^2\]

OpenStudy (anonymous):

Take the derivative. e*x^(e-1) -2x Set equal to 0 and solve. e*x^(e-1) = 2x x^(e-1) = 2/e *x x^(e-1) / x^1 = 2/e x^(e-1-1) = 2/e x^(e-2) = 2/e (x^(e-2))^(1/e-2) = (2/e)^(1/e-2) x = (2/e)^(1/e-2) which is about 1.65007509 Check f(x) at that critical point as well as at the endpoints, 0 and 2.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

My pleasure =D